Express the function 𝑓(π‘₯) = π‘₯ (π‘₯βˆ’3)2 as partial fractions and hence find ∫ 𝑓(π‘₯)𝑑�


1
Expert's answer
2021-06-10T07:05:51-0400

"f(x)=\frac{x}{(x-3)^2}"


"\frac{x}{(x-3)^2}=\frac{x-3+3}{(x-3)^2}=\frac{1}{x-3}+\frac{3}{(x-3)^2}"


"\int(\frac{1}{x-3}+\frac{3}{(x-3)^2})dx=log(x-3)-\frac{3}{x-3}+C"